OpenStudy (arti12):

prove that the volume of the largest cone that is inscribed in a sphere of radius R is 8/27 of the volume of sphere

OpenStudy (dumbcow):

 the diagram shows a relationship of: \[r^2 + (h-R)^2 = R^2\] solve for "h" to plug into Volume equation... \[h = \sqrt{R^2 - r^2} + R\] \[V = \frac{\pi}{3} r^2 h\] \[V = \frac{\pi}{3} r^2 (\sqrt{R^2 - r^2}+R)\] Next you want to maximize volume of cone, set derivative equal to 0 keeping "R" as a constant --> derivative using product rule: \[\frac{dV}{dr} = \frac{2\pi}{3} r (\sqrt{R^2 - r^2} +R) + \frac{\pi}{3} r^2(\frac{-r}{\sqrt{R^2 -r^2}}) = 0\] solve for "r^2" in terms of R ... \[r^2 = \frac{8R^2}{9}\] plug into Volume equation \[V = \frac{32}{81} \pi R^3\] Finally divide by volume of sphere to get ratio... \[\frac{(32/81) \pi R^3}{(4/3) \pi R^3} = \frac{8}{27}\]